Grade 11General Physics

a force F1 of 500 newton is required to push a car of mass 1000 kg slowly at a constant speed on a level road If force F2 of 1000 newton is applied the acceleration of the car will be

Aniket
8 Years agoGrade 11

2 Answers

Diwankar singh
8 Years ago
In first case:Force=500NMass=1000kg Therefore, F=m*a So, a=F/m a=500/1000 a=0.5m^2 In second case: Force=1000N Mass=1000kg a=F/m a=1000/1000 a=1m^2
Sahil
7 Years ago
Sice 500N of force is required to move the car with cosntant speed so out of 1000N , 500N of force will lost in moving the car with constant speed and the remaining 500N will provide acceleration . So
By the equation -- F=ma 
                             500=1000a 
                              a = 0.5 m/s